When a container is filled with 3.00 mol of H2, 2.00 mol of O2, and 1.00 mol of N2, the pressure in the container is 465 kpa. the partial pressure of O2 is...
1
Expert's answer
2017-06-19T06:14:33-0400
Solution:
x(O2)=2/(3+2+1)=0.333
p(O2)=P*x(O¬¬2)=465*0.333=155 (kpa)
Answer: The partial pressure of O2 is 155 kpa